Let X1,X2 and X3 be random variables with equal variances but with correlation coefficients ρ12=0.3,ρ13=0.5,ρ23=0.2 find the correlation coefficient of the linear function Y=X1+X2 and Y=X2+X3

1
Expert's answer
2021-10-13T15:08:39-0400

the correlation coefficient of the linear function Y=X1+X2 and Y=X2+X3:


"r_{X_1+X_2,X_2+X_3}=\frac{0.3\sigma^2+0.5\sigma^2+0.2\sigma^2}{\sigma_{X_1+X_2}\sigma_{X_2+X_3}}"


"Var(X_1+X_2)=Var(X_1)+2cov(X_1,X_2)+Var(X_2)=\sigma^2+2\cdot0.3\sigma^2+\sigma^2="

"=2.6\sigma^2"


similarly:

"Var(X_2+X_3)=\sigma^2+2\cdot0.2\sigma^2+\sigma^2=2.4\sigma^2"


"r_{X_1+X_2,X_2+X_3}=\frac{0.3\sigma^2+0.5\sigma^2+0.2\sigma^2}{\sqrt{2.6\sigma^2\cdot2.4\sigma^2}}=\frac{1}{\sqrt{2.6\cdot2.4}}=0.4003"